OpenStudy (anonymous):

I am confused to how I did this one. 7x>1/5. and the other one is 2 + 4x<18

sam (.sam.):

\[7x >\frac{1}{5}\] \[x >\frac{1}{5} \div7\] \[x >\frac{1}{5} \times\frac{1}{7}\] \[x >\frac{1}{35}\]

sam (.sam.):

2+4x<18 4x<16 x<4